Sally travels 450m west 950m south find the displacement and distance?
1
Expert's answer
2020-03-11T14:18:21-0400

The distance


"d=450+950=1400\:\rm m."

The displacement


"s=\sqrt{450^2+950^2}=1051\:\rm m."